# An Evaluation of Borda Count Variations Using Ranked Choice Voting Data

### v1

@@ 5 Analysis of Selected Elections
Table 5 displays the point totals for each of the five variations. Compared to the IRV outcome, the BC methods have much more fair results for this election. First, recall that monotonicity and no-show paradoxes cannot occur with these points-based methods. We also observe that Begich receives the most points with any method, which means there was no Condorcet winner failure, implying no verifiable failures are observed.
@@
Since Vance is a majority candidate with only two other candidates, the spoiler effect and compromise failures are easy to observe or dismiss. Removing Bryant from the slate would precipitate Vance now becoming the winner of the two-candidate election with ABC and MBC, resulting in a spoiler effect. Compromise failures, on the other hand, cannot occur with any of the five variations since there are such few Bryant$>$Flora$>$Vance or Bryant$>$Vance$>$Flora voters who could attempt this manipulation. For instance, the closest margin of victory is Flora winning over Vance using ABC by only 92.5 points. However, there are only 41 voters who could move Vance ahead of Bryant, leaving Vance still trailing Flora by 51.5 points.
@@
The next election we examine is the 2021 Minneapolis City Council Ward 2 Election, which featured five candidates - Tom Anderson, Yusra Arab, Guy Gaskin, Cam Gordon, and Robin Worlobah - where voters were only allowed to vote for up to three candidates in a truncated ballot. Note with this election, and some of the subsequent examples, we omit the preference profile because of its size. Despite only ranking up to three candidates, voters submitted 84 unique ballots that would need to be displayed. This election is unique in that it is one of only two elections without a Condorcet candidate. This is due to a Condorcet cycle (McCune and McCune 2023) in which Arab is preferred over Gordon, Gordon over Worlobah, and Worlobah over Arab. Worlobah ultimately won the election by less 20 votes in the final IRV round over Arab, while also leading in the initial plurality round by less than 50 votes. Our BC methods all agree on Arab as the winner, but since there is no majority or Condorcet candidate, there is no verifiable failure for either IRV or the BC variations. IRV was shown (Graham-Squire and McCune 2022), however, to have demonstrated both upward and downward monotonicity paradoxes, the spoiler effect, and a compromise failure.
@@
The 2022 Oakland, CA School Board District 4 election is very similar in that it also includes a Condorcet cycle between the three candidates Resnick, Hutchinson, and Manigo. Four of our BC methods agree on Resnick as the winner of this election, while MBC chooses the winner to be Hutchinson, who also was the IRV election winner. While Resnick has a large lead in first-place votes, as gleaned from Table 7, he also is listed third or left off of the most ballots, making this a very close election in terms of total points. Therefore, akin the Minneapolis election, this school board election features numerous manipulation failures with all five BC methods exhibiting compromise failures and spoiler effects, although only ABC and BCU are vulnerable to a truncation failure.
@@
The 2008 Pierce County, WA County Executive election, which featured a compromise failure in the IRV election (Graham-Squire and McCune 2022), is another example in which multiple failures exist for our BC methods. We first observe in this four-candidate election consisting of Shawn Bunney, Calvin Goings, Mike Lonergan, and Pat McCarthy, the five Borda variations do not choose the same winner. ABC, BCU, and MBC agree with IRV as having Pat McCarthy win the election, while EBC and QBC select Shawn Bunney, who we note is the plurality winner. Since McCarthy is the Condorcet winner, EBC and QBC exhibit a Condorcet winner failure. With this failure comes a spoiler effect for those two methods, although not for the other three, since removing Goings and Lonergan would result in McCarthy winning head-to-head.
@@
The 2009 Burlington, VT Mayor election is an election where the IRV included several unfair aspects, namely a monotonicity paradox, Condorcet winner failure, spoiler effect, and compromise failure (Graham-Squire and McCune 2022). Of the five candidates, the Condorcet winner was Montroll, while the IRV winner was Kiss. The five BC variations successfully select Montroll as the winner, avoiding any Condorcet failures as well as spoiler effects. There are manipulation failures present, including a compromise failure for EBC and QBC, as well as a truncation failure for EBC and BCU. For instance, when using BCU, Montroll defeats Kiss with 26194 points to 23427. If 692 of the 1328 voters whose top choice is Kiss followed in second by Montroll truncate their ballot to only list Kiss, then Montroll loses 4 points per ballots. This causes the winner to change from Montroll to Kiss, resulting in the truncation failure.
@@
The election where the BC methods exhibit the largest number of voting failures is the 2021 New York City Queens Borough President Democratic Primary election, seen in Table 8. In this three-candidate election with 195096 voters, the top two candidates Richards and Crowley were separated by only 192 first-place votes. The points for these two candidates across the five variations differed by as few as 432 points (344755 to 344323) for BCU and by at most 1597.5 points (502979.5 to 501382) for EBC and QBC. Not only do all five methods have a Condorcet winner failure by selecting Crowley over Richards, but these extremely close results also cause all five methods to have a spoiler effect and compromise failure, with only MBC avoiding a truncation failure. Recalling that MBC cannot exhibit this failure, we observe that for an election without a majority winner, every possible voting failure occurs within this single election.
@@ 6 Empirical Results
Our analysis of 421 U.S. RCV elections from 2004 to 2023 is displayed in Table 10, which shows the percentage of voting failure occurrences for each BC variation. Some rates use the complete set of 421 elections, but exceptions include the majority winner and loser failure rates, which are taken out of the 228 elections with a majority winner and the 109 with a majority loser. Likewise, the Condorcet winner rates use the 419 elections that had a Condorcet winner, and BCU’s Condorcet loser rate is based on the 420 elections that had such a candidate (only the previously discussed Oakland election did not). Finally, the spoiler effect rates use the 396 elections with 10 or fewer candidates.
@@
When comparing the manipulation failure rates, the BCU percentage for truncation failures illustrates the obvious incentive voters have to submit ballots that only include their top choice, known as bullet voting. Nearly half of the elections using this variation are vulnerable to truncation failures, with the three variations featuring the averaged approach for partial ballots performing much better. This occurs since truncating a ballot in those cases is a zero-sum strategy, meaning that your favored candidate will gain in points compared to other highly ranked candidates, but will instead lose ground over the lowest ranked candidates. BCU would instead cause your top choice to gain an advantage of all candidates who are removed from the ballot. The lower truncation failure rate for ABC compared to BCU mirrors the computation results by Kamwa (2022), which were referred to in that paper as the averaged and pessimistic models, respectively.
@@ 7 Conclusion
Further research is needed for these BC variations and for RCV methods in general on their outcomes with different types of elections. For example, do certain methods exhibit more voting failures when the number of candidates is higher, or within nonpartisan elections? Jungle primaries, which often lead to general elections with multiple candidates from the same party, could also lead to a higher risk of unfair outcomes. A deeper understanding of these dynamics could inform the design and adoption of monotonic voting methods such as variations of the Borda count.

### v2

@@ 5 Analysis of Selected Elections
Table 5 displays the point totals for each of the five variations. Compared to the IRV outcome, the BC methods have much more fair results for this election. First, recall that monotonicity and no-show paradoxes cannot occur with these points-based methods as they do with IRV. We also observe that Begich receives the most points with any method, which means there was no Condorcet winner failure, implying no verifiable failures are observed.
@@
Since Vance is a majority candidate with only two other candidates, the spoiler effect and compromise failures are easy to observe or dismiss. Removing Bryant from the slate would precipitate Vance now becoming the winner of the two-candidate election with ABC and MBC, resulting in a spoiler effect. Compromise failures, on the other hand, cannot occur with any of the five variations since there are such few Bryant$>$Flora$>$Vance or Bryant$>$Vance$>$Flora voters who could attempt this manipulation. For instance, the closest margin of victory is Flora winning over Vance using ABC by 92.5 points. However, there are only 41 voters who could move Vance ahead of Bryant, leaving Vance still trailing Flora by 51.5 points.
@@
The next election we examine is the 2021 Minneapolis City Council Ward 2 Election, which featured five candidates - Tom Anderson, Yusra Arab, Guy Gaskin, Cam Gordon, and Robin Worlobah - where voters were only allowed to vote for up to three candidates in a truncated ballot. Note with this election, and some of the subsequent examples, we omit the preference profile because of its size. Despite only ranking up to three candidates, voters submitted 84 unique ballots that would need to be displayed. This election is unique in that it is one of only two elections without a Condorcet candidate (McCune and McCune 2023). This is due to a Condorcet cycle in which Arab is preferred over Gordon, Gordon over Worlobah, and Worlobah over Arab. Worlobah ultimately won the election by less 20 votes in the final IRV round over Arab, while also leading in the initial plurality round by less than 50 votes. Our BC methods all agree on Arab as the winner, but since there is no majority or Condorcet candidate, there is no verifiable failure for either IRV or the BC variations. IRV was shown (Graham-Squire and McCune 2022), however, to have demonstrated both upward and downward monotonicity paradoxes, the spoiler effect, and a compromise failure.
@@
The 2022 Oakland, CA School Board District 4 Election is very similar in that it also includes a Condorcet cycle between the three candidates Resnick, Hutchinson, and Manigo. Four of our BC variations agree on Resnick as the winner of this election, while MBC chooses the winner to be Hutchinson, who also was the IRV election winner. While Resnick has a large lead in first-place votes, as gleaned from Table 7, he is also listed third or left off of the most ballots, making this a very close election in terms of total points. Therefore, akin the Minneapolis election, this school board election features numerous manipulation failures with all five BC methods exhibiting compromise failures and spoiler effects, although only ABC and BCU are vulnerable to a truncation failure.
@@
The 2008 Pierce County, WA County Executive Election, which featured a compromise failure in the IRV election (Graham-Squire and McCune 2022), is another example in which multiple failures exist for our BC methods. We first observe in this four-candidate election consisting of Bunney, Goings, Lonergan, and McCarthy, the five Borda variations do not choose the same winner. ABC, BCU, and MBC agree with IRV as having Pat McCarthy win the election, while EBC and QBC select Shawn Bunney, who we note is the plurality winner. Since McCarthy is the Condorcet winner, EBC and QBC exhibit a Condorcet winner failure. With this failure comes a spoiler effect for those two methods, although not for the other three, since removing Goings and Lonergan would result in McCarthy winning head-to-head.
@@
The 2009 Burlington, VT Mayor Election is an election where the IRV result included several unfair aspects, namely a monotonicity paradox, Condorcet winner failure, spoiler effect, and compromise failure (Graham-Squire and McCune 2022). Of the five candidates, the Condorcet winner was Montroll, while the IRV winner was Kiss. The five BC variations successfully select Montroll as the winner, avoiding any Condorcet failures as well as spoiler effects. There are manipulation failures present, including a compromise failure for EBC and QBC, as well as a truncation failure for EBC and BCU. For instance, when using BCU, Montroll defeats Kiss with 26194 points to 23427. If 692 of the 1328 voters whose top choice is Kiss followed in second by Montroll truncate their ballot to only list Kiss, then Montroll loses 4 points per ballots. This causes the winner to change from Montroll to Kiss, resulting in the truncation failure.
@@
The election where the BC variatiosn exhibit the largest number of voting failures is the 2021 New York City Queens Borough President Democratic Primary Election, seen in Table 8. In this three-candidate election with 195096 voters, the top two candidates, Richards and Crowley, were separated by only 192 first-place votes. The points for these two candidates across the five variations differed by as few as 432 points (344755 to 344323) for BCU and by at most 1597.5 points (502979.5 to 501382) for EBC and QBC. Not only do all five methods have a Condorcet winner failure by selecting Crowley over Richards, but these extremely close results also cause all five methods to have a spoiler effect and compromise failure, with only MBC avoiding a truncation failure. Recalling that MBC cannot exhibit this failure, we observe that for an election without a majority winner, the maximum number of voting failures occur within this single election.
@@ 6 Empirical Results
Our analysis of 421 U.S. RCV elections from 2004 to 2023 is displayed in Table 10, which shows the percentage of voting failure occurrences for each BC variation. Some rates use the complete set of 421 elections, but exceptions include the majority winner and loser failure rates, which are taken out of the 228 elections with a majority winner and the 109 with a majority loser. Likewise, the Condorcet winner rates use the 419 elections that had a Condorcet winner, and BCU’s Condorcet loser rate is based on the 420 elections that had such a candidate (only the Oakland school board election did not). Finally, the spoiler effect rates use the 396 elections with 10 or fewer candidates.
@@
When comparing the manipulation failure rates, the BCU percentage for truncation failures illustrates the obvious incentive voters have with this method to submit ballots that only include their top choice, known as bullet voting. Nearly half of the elections using this variation are vulnerable to truncation failures, with the three variations featuring the averaged approach for partial ballots performing much better. This occurs since truncating a ballot in those cases is a zero-sum strategy, meaning that your favored candidate will gain in points compared to other highly ranked candidates, but will instead lose ground over the lowest ranked candidates. BCU would instead cause your top choice to gain an advantage over all candidates who are removed from the ballot. The lower truncation failure rate for ABC compared to BCU mirrors the computation results by Kamwa (2022), which were referred to in that paper as the averaged and pessimistic models, respectively.
@@ 7 Conclusion
More research is also needed for these BC variations and for RCV methods in general on their outcomes with different types of elections. For example, do certain methods exhibit more voting failures when the number of candidates is higher, or within nonpartisan elections? Jungle primaries, which often lead to general elections with multiple candidates from the same party, could also lead to a higher risk of unfair outcomes. A deeper understanding of these dynamics could inform the design and adoption of monotonic voting methods such as variations of the Borda count.